OpenStudy (anonymous):

George rides his bike at 20 miles per hour. He travels 5 miles. How many minutes will this take him?

OpenStudy (anonymous):

distance = velocity * time so, time = distance/velocity Just change the 20mi/hr into 20 mi/60 minutes and your answer will be in minutes.

OpenStudy (anonymous):

ohhh okay i was using a wrong formula haha thanks. so its 30

OpenStudy (anonymous):

I'm getting 15, doing it in my head.

OpenStudy (anonymous):

t = 5 miles/ (20/60) = 5 * 3 = 15

OpenStudy (anonymous):

oh i get it now. thank you! :)

OpenStudy (anonymous):

Welcome